Suppose that X has a distribution N(10, 1). Determine c such that P(X<c)=2P(X>c)


1
Expert's answer
2021-11-18T14:57:03-0500
"P(X<c)=2P(X>c)=2(1-P(X\leq c))"

"P(X<c)=2-2P(X\leq c)"

"P(X<c)=\dfrac{2}{3}"

"P(Z<\dfrac{c-10}{1})=\dfrac{2}{3}"

"c-10=0.4307273"

"c=10.4307273"